OpenStudy (anonymous):

if A & B are second quadrant angles such that sinA=2/3 & cosB=-1/2, find (a) sin(A+B) (b) cos(A+B) (c) tan(A+B)

OpenStudy (anonymous):

ok what you need are the following numbers: \(\cos(A)\) and \(\sin(B)\) so you can use the various formulas do you know how to get them

OpenStudy (anonymous):

uhh you would use pythagorean ?

OpenStudy (anonymous):

for \(\sin(A)=\frac{2}{3}\) i would for \(\cos(B)=\frac{1}{2}\) i would know instantly that \(\sin(B)=\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

\[\cos(A) = \frac{ \sqrt{5} }{ 3}\]

OpenStudy (anonymous):

then i would use the equation\[\sin(u+v)=\sin u \cos v + \cos u \sin v\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait... if angle B is a second quadrant angle, then shouldn't cos(B) = negative?

OpenStudy (anonymous):

@Byteme good point !

OpenStudy (anonymous):

oh ... sorry i meant it to be -1/2

OpenStudy (anonymous):

oh ok good still the same sine though but \(\cos(A)=-\frac{\sqrt5}{3}\)

OpenStudy (anonymous):

ye. so id substitute each to the sum & difference formulas?

OpenStudy (anonymous):

yep...